# Scalable Causal Structure Learning via Amortized Conditional Independence Testing

### v2

@@ 4 Conclusion
We introduced a new algorithm for causal discovery which drastically decreases the computational burden required to compute a $p$-value for a causal relationship between two nodes in a directed acyclic graph with temporally separated sets of variables. We tested this methodology on semi-synthetic data constructed from a recent study on somatic tumor mutations and metastatic potential for a panel of patients and found that the methodology successfully controlled Type I error and had reasonable power across datasets of differing levels of confoundedness. When run on the dataset of Nguyen et al. [2022] , interesting connections between metastases and genes are identified.
Several avenues for follow-up work exist. From a statistical perspective, the $p$-values generated from the procedure are conservative and power can potentially be improved through post-hoc adjustment to the $p$-values, for example through the use of Empirical Bayes methods [Efron, 2008]. From a computational perspective, more sophisticated probability models could be used to find the worst-case $p$-value to search the combinatorial space.
@@ Appendix B Algorithms for construction of semi-synthetic datasets
     For each elements $Y_{k}\in\mathcal{S}_{2}$ with $l<k$, add $Y_{l}$ to $\mathcal{Y}_{Y_{k}}^{\star}$ with probability $p$.
@@ Appendix C Adjustments for continuous and mixed-value data
In Algorithm 4, we present a slightly augmented the methodology to train amortized predictive models to accommodate continuous valued data. The algorithm is largely the same as the one presented in the main paper, with the modification that both the masks $(B,M)$ and the interactions between the mask are inputs into the loss function. In the case that binary-valued data was used with $X_{j},Y_{k}\in\{-1,1\}$, then this was not necessary as the interactions would simply lead to the masked versions $\tilde{X}_{j},\tilde{Y}_{k}\in\{-1,0,1\}$. In the case of continuous or mixed valued data, however, there is a need to distinguish between data that is 00 because it has been masked or data that is actually 00.
@@ Appendix E Technical exposition of Generalized Covariance Measure
We recall the details of the technical conditions described in Shah and Peters [2020] for completeness. We assume that the dataset consists of i.i.d. $n$ samples, with individual observations denoted $\{X_{j}^{i},Y_{k}^{i},S^{i},X_{-j}^{i}\}_{i=1}^{n}$. Let $\mathcal{P}_{0}$ denote the family of joint distributions corresponding to the null that $X_{j}\perp\!\!\!\perp Y_{k}|S,X_{-j}$.
@@
Fact E.1 (Theorem 6 of Shah and Peters [2020]).
@@ Appendix F Additional experimental results
Although the PC-p algorithm is the only existing method that aims at constructing $p$-values with frequentist error control, we also benchmark this procedure against approaches not aimed at controlling type I error that are implemented in the TETRAD project. The only two methods that scale reasonably to datasets of a similar size as the dataset of Nguyen et al. [2022] are Fast Greedy Equivalence Search (FGES) [Ramsey et al., 2017b] and the Best Order Score Search (BOSS) algorithm [Lam et al., 2022]. Other methods used as baselines are Cyclic Causal Discovery (CCD, Richardson [1996]) , Fast Causal Inference (FCI, Spirtes [2001]), Greedy Fast Causal Inference (GFCI, Ogarrio et al. [2016]), Greedy relaxation of the sparsest permutation (GRaSP, Lam et al. [2022]), and GRaSP-FCI [Ogarrio et al., 2016].

### v3

@@ 4 Conclusion
We introduced a new algorithm for causal discovery which drastically decreases the computational burden required to compute a $p$-value for a causal relationship between two nodes in a directed acyclic graph with temporally separated sets of variables. We tested this methodology on semi-synthetic data constructed from a recent study on somatic tumor mutations and metastatic potential for a panel of patients and found that the methodology successfully controlled Type I error and had reasonable power across datasets of differing levels of confoundedness. When run on the dataset of data_source , interesting connections between metastases and genes are identified.
Several avenues for follow-up work exist. From a statistical perspective, the $p$-values generated from the procedure are conservative and power can potentially be improved through post-hoc adjustment to the $p$-values, for example through the use of Empirical Bayes methods [efron]. From a computational perspective, more sophisticated probability models could be used to find the worst-case $p$-value to search the combinatorial space.
@@ Appendix B Algorithms for construction of semi-synthetic datasets
  For each elements $Y_{k}\in\mathcal{S}_{2}$ with $l<k$, add $Y_{l}$ to $\mathcal{Y}_{Y_{k}}^{\star}$ with probability pp.
@@ Appendix C Adjustments for continuous and mixed-value data
In Algorithm 4, we present a slightly augmented the methodology to train amortized predictive models to accommodate continuous valued data. The algorithm is largely the same as the one presented in the main paper, with the modification that both the masks $(B,M)$ and the interactions between the mask are inputs into the loss function. In the case that binary-valued data was used with $X_{j},Y_{k}\in\{-1,1\}$, then this was not necessary as the interactions would simply lead to the masked versions $\tilde{X}_{j},\tilde{Y}_{k}\in\{-1,0,1\}$. In the case of continuous or mixed valued data, however, there is a need to distinguish between data that is 0 because it has been masked or data that is actually 0.
@@ Appendix E Technical exposition of Generalized Covariance Measure
We recall the details of the technical conditions described in hardness_CI for completeness. We assume that the dataset consists of i.i.d. $n$ samples, with individual observations denoted $\{X_{j}^{i},Y_{k}^{i},S^{i},X_{-j}^{i}\}_{i=1}^{n}$. Let $\mathcal{P}_{0}$ denote the family of joint distributions corresponding to the null that $X_{j}\perp\!\!\!\perp Y_{k}|S,X_{-j}$.
@@
Fact E.1 (Theorem 6 of hardness_CI).
@@ Appendix F Additional experimental results
Although the PC-p algorithm is the only existing method that aims at constructing $p$-values with frequentist error control, we also benchmark this procedure against approaches not aimed at controlling type I error that are implemented in the TETRAD project. The only two methods that scale reasonably to datasets of a similar size as the dataset of data_source are Fast Greedy Equivalence Search (FGES) [fges] and the Best Order Score Search (BOSS) algorithm [greedy_sparest]. Other methods used as baselines are Cyclic Causal Discovery (CCD, CCD) , Fast Causal Inference (FCI, FCI), Greedy Fast Causal Inference (GFCI, gfci), Greedy relaxation of the sparsest permutation (GRaSP, greedy_sparest), and GRaSP-FCI [gfci].